OpenStudy (anonymous):

What is the area of a sector with radius 10" and measure of arc equal to 45°? 2.5 pi sq. in. 12.5 pi sq. in. 25 pi sq. in.

OpenStudy (anonymous):

whats the radius can't see

OpenStudy (nurali):

a = A(45)/360 a = (pir^2)(1/8) a = pi(10)^2(1/8) a = 12.5 pi sq. in.

OpenStudy (anonymous):

wow r u coming here for helping others or getting medals....no offense